OpenStudy (anonymous):

\[f(x)=\frac{x}{2x^2 +1}=\] \[x*\frac{1}{2x^2+1}\] I don't know how to write this as a sum

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}\]

OpenStudy (anonymous):

n is varying or x is varying?

OpenStudy (anonymous):

I'm not sure, It's a power series

OpenStudy (anonymous):

here is an example from my book \[\frac{x^3}{(x+2)}= x^3*\frac{1}{x+2}=x^3\sum_{n=0}^{\infty}\frac{(-1)^n}{2^{n+1}}x^n=\sum_{n=0}^{\infty}\frac{(-1)^n}{2^{n+1}}x^{n+3} \]

OpenStudy (anonymous):

firstly tell me do you know what is a power series?

OpenStudy (anonymous):

here is what I under stand... we have a function like \[\frac{x^3}{(x+2)}\] and we are trying to express it as a sum of infinitely many terms. so we're trying to trace the line : \[\frac{x^3}{x+2}\] with partial sums...

OpenStudy (anonymous):

I think something just clicked

OpenStudy (anonymous):

\[\sum c_n(x-a)^n=c_0 +c_1(x-a)^2+....\]

OpenStudy (anonymous):

\[x=c_n\] \[(x-a)=\]

OpenStudy (anonymous):

a=-1

OpenStudy (anonymous):

check them they will help you

OpenStudy (anonymous):

\[x=2x^2\]?

OpenStudy (helder_edwin):

Are you looking for the Taylor series of the function you wrote?

OpenStudy (anonymous):

no, Power series

OpenStudy (anonymous):

\[\sum c_n(x-a)^n=c_0 +c_1(x-a)^2+....\]

OpenStudy (helder_edwin):

you gave an example from a book. which book is it?

OpenStudy (anonymous):

stewart

OpenStudy (anonymous):

6th edition I believe

OpenStudy (helder_edwin):

calculus?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I think \[c_n =x\] in this problem

OpenStudy (anonymous):

would that make any sense? I'm new to power series

OpenStudy (anonymous):

I can write the whole theorem if you'd like

OpenStudy (helder_edwin):

yes that might help

OpenStudy (helder_edwin):

is it one-variable calculus or multivariable? the book.

OpenStudy (anonymous):

early transcendentals

OpenStudy (anonymous):

i gues that's one variable

OpenStudy (helder_edwin):

does it have violins on the cover?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If the power series \[\sum c_n(x-a)^n\] has a radius of convergence R>0 then the function f defined by \[f(x)=c_0 +c_1(x-a)+c_2(x-a)^2+....=\sum_{n=0}^{\infty}c_n(x-a)^n\] is differentiable ( and therefore continuous) on the interval (a-R, a+R) and (i) \[f'(x)=c_1+2c_2(x-a)+3c_3(x-a)^2+....=\sum_{n=1}^{\infty} nc_n(x-a)^{n-1}\] (ii) \[\int f(x)dx=C+c_0(x-a)+c_1\frac{(x-a)^2}{2}....=C+\sum{n=0}^{\infty}c_n\frac{(x-a)^{n+1}}{n+1}\]

OpenStudy (anonymous):

the radii of convergence of the power series in equations (i) and (ii) are both R

OpenStudy (anonymous):

that was quite tedious

OpenStudy (helder_edwin):

i bet

OpenStudy (helder_edwin):

so. are you looking for a series that converges to \[ \frac{x}{2x^2+1} \] i.e. \[ \frac{x}{2x^2+1}=\sum_{n=0}^\infty c_n(x-a)^n \]

OpenStudy (anonymous):

the question reads Find a power series representation for the function and determine the interval of convergence

OpenStudy (helder_edwin):

give a minute ok?

OpenStudy (anonymous):

sure, I've been trying to figure this out for 2 days now

OpenStudy (helder_edwin):

does it give a point for the convergence? i mean, you can't have a power series to cenverge to the function "everywhere"

OpenStudy (anonymous):

In the example that I gave, the interval of convergence is (-2,2)

OpenStudy (anonymous):

\[\frac{x^3}{(x+2)}= x^3*\frac{1}{x+2}=x^3\sum_{n=0}^{\infty}\frac{(-1)^n}{2^{n+1}}x^n=\sum_{n=0}^{\infty}\frac{(-1)^n}{2^{n+1}}x^{n+3}\]

OpenStudy (anonymous):

I will be right back

OpenStudy (helder_edwin):

are you back?

OpenStudy (helder_edwin):

an

OpenStudy (helder_edwin):

anyway

OpenStudy (helder_edwin):

in the example you just gave me if \[ f(x)=\frac{1}{x+2} \] then \[ f^{(n)}=\frac{(-1)^{n+1}\cdot n!}{(x+2)^{n+1}} \] do you agree?

OpenStudy (anonymous):

I'm back

OpenStudy (anonymous):

not quite...why is it n+1?

OpenStudy (anonymous):

is it because it's even

OpenStudy (helder_edwin):

you are right. it should be only n

OpenStudy (anonymous):

well the book has \[(x+2)^{n+1}\] why is it n+1?

OpenStudy (helder_edwin):

ok

OpenStudy (helder_edwin):

the consecutive derivatives would be \[ f'(x)=\frac{-1}{(x+2)^2} \] \[ f''(x)=\frac{2}{(x+2)^3} \] \[ f'''(x)=\frac{-6}{(x+2)^4} \] do you agree?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ooh hence n+1?

OpenStudy (helder_edwin):

so the n-th derivative is \[ f^{(n)}(x)=\frac{(-1)^n\cdot n!}{(x+2)^{n+1}} \] agree?

OpenStudy (anonymous):

yes sir! one more thing...why n! ?

OpenStudy (helder_edwin):

notice de numerators of the fractions. they are 1=1!, 2=2! and 6=3! ok?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

got it

OpenStudy (helder_edwin):

for instance \[ f^{(4)}(x)=\frac{24}{(x+2)^5} \]

OpenStudy (anonymous):